NEXT
Make New Entry,
Make Followup Entry
User name R. Subedi
Log entry time 08:23:31 on January19,2009
Entry number 256790
keyword=Run Coordinator Report: Jan 18, 2009
Run Coordinator Report for Jan 18, 2009:
# Polarity status of the spectrometers: BigBite -ve (for e-), L-HRS +ve (for &pi+)
# &pi+Transverse mode before 10:00 with the half-wave-plate IN
# &pi+Vertical mode after 14:00 with the half-wave-plate IN
Charge Accumulated Summary So Far
Half Wave Plate Charge Expected (C) Charge Accumulated (C) Comment
&pi- Vertical IN 3.5 0.86 -
&pi- Vertical OUT 3.5 4.04 Done
&pi-Transverse IN 3.5 0.62 -
&pi- Transverse OUT 3.5 4.38 Done
&pi+ Vertical IN 3.5 2.56 Currently here
&pi+ Vertical OUT 3.5 3.72 Done
&pi+ Transverse IN 3.5 4.5 Done
&pi+ Transverse OUT 3.5 3.66 Done
Overall Plans
1. Will stay at Pi-plus Vertical mode until Thursday Morning.
2. Will do Moller measurement either on Tuesday or Wednesday.
3. Will do positron runs on Thursday (Jan 22) Day/Swing and Swing/Owl(Friday) where
Day/Swing [total one shift, starting ~11:00]: BigBite +ve polarity and L-HRS +ve
Swing/Owl [total one shift, starting ~19:00]: BigBite +ve polarity and L-HRS -ve
4. Friday (Jan 23) Day: BigBite will be back to -ve polarity, L-HRS stays at -ve polarity.
Activities
- Production in &pi+ Transverse mode until 10:00.
- 10:00 starts target polarization change from Transverse to Vertical.
Also asked for opportunistic controlled access to do misc DAQ work
(L-HRS, RICH) in the hall.
- 13:50 Target work (Transverse to Vertical polarization change) finished,
asked for beam tuning for a Compton-quality.
- 14:30 Reference cell (H2) data taking starts
- 18:00 Production on polarizide 3He on TRANSVERSE mode starts and continues smoothly.
Beamtime Accounting:
ABU BANU BNA Comment
Owl 6.87 0.17 0.98 half an hour beam stop for MCC work
Day 2.65 4.36 1.00 Target polarztion state change (T->V), beam tuning.
Swing 6.61 1.14 0.28